[xv Yv (x)]= xv Yv-1(x)


1
Expert's answer
2021-03-15T08:20:13-0400

To prove:

"{d\over dx}[x^vY_{v}(x)]=x^vY_{v-1}(x)"


Taking Left Hand Side:

"{d\over dx}[x^vY_{v}(x)]"


We have to assume "m=Y_u(x)\ \&\ n=x^v"

Now, use Leibnitz Theorem

"{d\over dx}[x^vY_{v}(x)]=^1C_0(1!\times x^v \times Y_{v-1}(x))"

"=" "x^vY_{v-1}(x)=" Right Hand Side


Hence "{d\over dx}[x^vY_{v}(x)]=x^vY_{v-1}(x)"